OpenStudy (anonymous):

In the electron cloud model, the electron cloud is denser in some locations than in others. What do the denser areas represent?

OpenStudy (anonymous):

A: regions where the probability of finding an electron is high B: orbitals that contain the greatest number of electrons C: energy levels that are the least stable D: electrons that have the highest energy

OpenStudy (aaronq):

"cloud model" i hate that term lol but the answer would be "regions where the probability of finding an electron is high"

OpenStudy (frostbite):

Would you rather write: \[\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}\psi ^{2} dx dy dz=1\] Or normalized: \[\int\limits_{-\infty}^{\infty}\psi ^{2}d \tau=1\]

OpenStudy (aaronq):

much much better. haha